News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 12 August 2022 9.00am EDT

During the last week the Chandra observing schedule was interrupted by an entry to Bright-Star Hold (BSH) at 9:49am EDT on Aug. 11. The entry into BSH was nominal and an assessment of the telemetry indicated that all spacecraft hardware was performing nominally. The cause of BSH entry was confirmed to be the high ionizing radiation environment near perigee, which interrupted the ability to track the desired guide stars.

Real-time procedures were executed on Aug 5-8 to dump OBC-A and -B memory as a follow-up to the flight software patch uplinked on Aug 4 to re-configure the on board software to power down the HRC 15 volt power supply during a safing action. The dumps will be used to update the baseline memory images maintained on the ground.

Further real-time procedures were executed on Aug 11 to update the ACIS Flight Software to version 58 (Standard G, Optional I) and to update the txings parameters to optimized values appropriate for the current state of the solar cycle. These actions will improve the ability of ACIS to serve as a monitor of the radiation environment and to safe the science instruments under conditions of elevated radiation.

Additional real-time procedures were executed on Aug 11 to initiate recovery from BSH. These procedures restarted aspect camera processing in the OBC, followed by a manually-commanded maneuver to a solar pitch angle of 156 deg to begin cooling and preserve a safe momentum profile prior to science resumption. As clean-up actions, real-time procedures were executed to disable SCS-29 and to re-enable SCS 107 in preparation for science resumption.

The next schedule, with a planned start of Aug 13, is still in work. This and future loads will mitigate the risk of entry into BSH by the use of gyro hold during passage through the elevated radiation region near perigee.

All spacecraft subsystems except HRC continued to support nominal operations.

Return to Status Reports